# Support Recovery and $\ell_{2}$-Error Bound for Sparse Regression with Quadratic Measurements via Weakly-Convex-Concave Regularization

### v1

@@ 0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ 1
Table 5: MSE(SD) of relative error for $d=128,s=10$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.091 (0.1414) 1.114 (0.1394) 1.091 (0.1646) 1.111 (0.1571) 1.152 (0.1743) 1.166 (0.1704) 1.187 (0.1779)
0.2 MSE 1.068 (0.1778) 1.108 (0.1568) 1.042 (0.1955) 1.03 (0.2445) 1.12 (0.3158) 1.133 (0.2951) 1.122 (0.349)
0.3 MSE 0.9705 (0.293) 1.052 (0.2643) 0.9269 (0.2784) 0.8665 (0.3783) 0.7603 (0.6056) 0.9416 (0.4998) 0.823 (0.5875)
0.4 MSE 0.5529 (0.5201) 0.7402 (0.4788) 0.6218 (0.407) 0.4724 (0.4664) 0.4097 (0.5788) 0.5031 (0.5934) 0.4627 (0.6022)
0.5 MSE 0.3079 (0.5235) 0.4849 (0.5386) 0.4257 (0.4788) 0.3526 (0.477) 0.2461 (0.4983) 0.2959 (0.5359) 0.2387 (0.4991)
0.6 MSE 0.05125 (0.2487) 0.2159 (0.4243) 0.2071 (0.3718) 0.06997 (0.2575) 0.04582 (0.2242) 0.08147 (0.2997) 0.04663 (0.2289)
0.7 MSE 0.07634 (0.302) 0.08627 (0.316) 0.1135 (0.3098) 0.05085 (0.2215) 0.07238 (0.2859) 0.07605 (0.3014) 0.05124 (0.2502)
0.8 MSE 0.01417 (0.1356) 0.008939 (0.08315) 0.02032 (0.141) 0.01137 (0.1108) 0.0005478 (0.0003627) 0.0003913 (0.0003058) 0.0004205 (0.0003104)
0.9 MSE 0.01311 (0.1256) 0.02341 (0.1631) 0.01126 (0.1101) 0.02173 (0.1514) 0.01205 (0.1157) 0.01312 (0.1279) 0.012 (0.1164)
1.0 MSE 0.0005213 (0.0002837) 0.0004813 (0.0008072) 0.008467 (0.08181) 0.0002522 (0.0001432) 0.0004765 (0.0002836) 0.0003329 (0.0002278) 0.0003577 (0.0002334)
Table 6: MSE(SD) of relative error for $d=128,s=10$ and $\sigma=0.05$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.085 (0.1328) 1.106 (0.1324) 1.087 (0.1423) 1.101 (0.1628) 1.153 (0.1718) 1.16 (0.1624) 1.175 (0.1777)
0.2 MSE 1.049 (0.1958) 1.101 (0.1642) 1.021 (0.2048) 1.043 (0.2243) 1.094 (0.3356) 1.121 (0.2632) 1.13 (0.342)
0.3 MSE 0.8273 (0.3511) 0.9234 (0.3213) 0.8148 (0.3337) 0.7896 (0.3815) 0.6387 (0.593) 0.6949 (0.5856) 0.6682 (0.612)
0.4 MSE 0.6436 (0.5085) 0.8282 (0.4484) 0.6899 (0.395) 0.5693 (0.4823) 0.4613 (0.5924) 0.59 (0.5957) 0.5024 (0.607)
0.5 MSE 0.243 (0.4811) 0.509 (0.5236) 0.4271 (0.471) 0.2844 (0.4427) 0.2429 (0.4918) 0.2886 (0.5111) 0.205 (0.4681)
0.6 MSE 0.203 (0.4607) 0.2593 (0.493) 0.2211 (0.405) 0.1493 (0.3698) 0.174 (0.4275) 0.1916 (0.454) 0.1882 (0.4472)
0.7 MSE 0.02853 (0.1732) 0.07949 (0.2845) 0.08671 (0.2745) 0.02283 (0.1482) 0.02904 (0.1793) 0.03935 (0.211) 0.04164 (0.2227)
0.8 MSE 0.003159 (0.002873) 0.03442 (0.1868) 0.05568 (0.2171) 0.02166 (0.1408) 0.002817 (0.002608) 0.0135 (0.1145) 0.01544 (0.1327)
0.9 MSE 0.0153 (0.1272) 0.01504 (0.1357) 0.01208 (0.1083) 0.01222 (0.1099) 0.01495 (0.1276) 0.01413 (0.1258) 0.01482 (0.1317)
1.0 MSE 0.002603 (0.001302) 0.0154 (0.1383) 0.01241 (0.1108) 0.01251 (0.112) 0.01539 (0.1304) 0.01415 (0.1254) 0.001726 (0.0009298)
Table 7: MSE(SD) of relative error for $d=256,s=15$ and $\sigma=0.01$
$n/d$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP Firm LOG EXP
0.1 MSE 1.121 (0.06421) 1.143 (0.05974) 1.138 (0.09982) 1.144 (0.07845) 1.18 (0.1058) 1.231 (0.08987) 1.263 (0.1073)
0.2 MSE 1.103 (0.15) 1.147 (0.1259) 1.062 (0.2094) 1.049 (0.2355) 1.087 (0.3919) 1.165 (0.3045) 1.15 (0.377)
0.3 MSE 0.9678 (0.3185) 1.101 (0.2053) 0.8716 (0.3311) 0.866 (0.3872) 0.7233 (0.6405) 0.8856 (0.5504) 0.7427 (0.6431)
0.4 MSE 0.5304 (0.5944) 0.8253 (0.482) 0.5715 (0.497) 0.4767 (0.5222) 0.3911 (0.5887) 0.5146 (0.6151) 0.4587 (0.6179)
0.5 MSE 0.2341 (0.476) 0.4294 (0.534) 0.293 (0.4405) 0.1878 (0.4069) 0.1399 (0.3996) 0.2263 (0.4728) 0.1496 (0.4082)
0.6 MSE 0.1276 (0.3851) 0.1752 (0.4071) 0.1024 (0.3097) 0.08888 (0.3035) 0.07372 (0.2932) 0.08987 (0.3296) 0.08984 (0.329)
0.7 MSE 0.06553 (0.2857) 0.08276 (0.3045) 0.04321 (0.2121) 0.04276 (0.21) 0.06264 (0.2731) 0.06505 (0.2842) 0.05078 (0.2491)
0.8 MSE 0.0003446 (0.000144) 0.02425 (0.1695) 0.01038 (0.1022) 0.02158 (0.1507) 0.0003082 (0.0001376) 0.0002127 (0.0001035) 0.0002275 (0.0001077)
0.9 MSE 0.0002978 (0.0001248) 0.0001692 (0.0001473) 0.01013 (0.09999) 0.0001356 (6.212e-05) 0.0002639 (0.0001202) 0.0001809 (9.108e-05) 0.0001931 (9.303e-05)
[…]

### v2

@@ 0 $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
Table 6: RelErr for $p=128$, $s=10$, $\sigma=10^{-2}$.
$n/p$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 1.10 (0.11) 1.11 (0.11) 1.12 (0.12) 1.11 (0.12) 1.20 (0.15) 1.20 (0.14) 1.17 (0.12)
0.2 1.06 (0.17) 1.08 (0.14) 1.06 (0.19) 1.05 (0.18) 1.10 (0.36) 1.13 (0.25) 1.15 (0.24)
0.3 8.40e-1 (3.65e-1) 9.54e-1 (2.93e-1) 8.14e-1 (3.80e-1) 8.45e-1 (3.29e-1) 6.39e-1 (6.16e-1) 7.46e-1 (5.65e-1) 8.04e-1 (5.59e-1)
0.4 4.93e-1 (5.40e-1) 7.02e-1 (4.76e-1) 4.60e-1 (4.89e-1) 4.91e-1 (4.68e-1) 3.45e-1 (5.63e-1) 4.12e-1 (5.59e-1) 4.40e-1 (5.98e-1)
0.5 2.30e-1 (4.80e-1) 3.88e-1 (5.37e-1) 2.49e-1 (4.66e-1) 2.34e-1 (4.43e-1) 1.66e-1 (4.33e-1) 2.42e-1 (4.87e-1) 2.70e-1 (5.18e-1)
0.6 9.54e-2 (3.25e-1) 2.11e-1 (4.33e-1) 9.74e-2 (3.13e-1) 1.14e-1 (3.29e-1) 7.39e-2 (2.95e-1) 8.59e-2 (3.06e-1) 1.10e-1 (3.54e-1)
0.7 1.20e-2 (1.18e-1) 8.37e-2 (2.89e-1) 1.81e-2 (1.27e-1) 3.52e-2 (1.74e-1) 1.21e-2 (1.20e-1) 3.50e-2 (1.92e-1) 1.18e-2 (1.17e-1)
0.8 1.40e-2 (1.39e-1) 2.38e-2 (1.68e-1) 1.23e-2 (1.23e-1) 2.21e-2 (1.55e-1) 1.29e-2 (1.29e-1) 2.49e-2 (1.74e-1) 2.55e-2 (1.79e-1)
0.9 1.19e-2 (1.18e-1) 1.25e-2 (1.23e-1) 9.82e-3 (9.77e-2) 1.01e-2 (1.00e-1) 7.39e-5 (3.12e-5) 1.12e-2 (1.11e-1) 1.25e-2 (1.24e-1)
1.0 9.33e-5 (3.14e-5) 3.07e-4 (1.06e-3) 4.89e-5 (1.72e-5) 4.60e-5 (1.70e-5) 7.16e-5 (2.60e-5) 1.97e-4 (1.04e-3) 6.04e-5 (2.24e-5)
Table 7: Computation time (seconds) for $p=128$, $s=10$, $\sigma=10^{-2}$.
$n/p$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 1.02e-2 (2.74e-3) 2.71e-2 (1.68e-2) 5.26e-3 (3.94e-3) 4.29e-3 (1.25e-3) 1.69e-1 (4.94e-2) 7.49e-2 (1.46e-2) 1.21e-1 (6.56e-2)
0.2 2.37e-2 (6.71e-3) 6.37e-2 (2.24e-2) 4.56e-3 (1.89e-3) 4.75e-3 (1.25e-3) 2.14e-1 (5.78e-2) 9.10e-2 (1.60e-2) 2.59e-1 (1.40e-1)
0.3 7.10e-2 (3.65e-2) 9.25e-2 (5.48e-2) 6.69e-3 (3.37e-3) 6.82e-3 (4.27e-3) 1.55e-1 (9.05e-2) 9.94e-2 (6.63e-2) 4.66e-1 (2.69e-1)
0.4 2.61e-1 (1.74e-2) 1.33e-1 (8.08e-2) 2.61e-2 (1.54e-2) 2.39e-2 (1.10e-2) 1.78e-1 (1.37e-1) 1.70e-1 (1.29e-1) 7.44e-1 (7.38e-1)
0.5 2.82e-1 (2.76e-2) 1.10e-1 (8.11e-2) 3.28e-2 (1.25e-2) 2.77e-2 (9.24e-3) 1.45e-1 (1.41e-1) 1.38e-1 (1.27e-1) 6.10e-1 (8.68e-1)
0.6 2.24e-1 (6.83e-2) 7.46e-2 (7.64e-2) 3.25e-2 (8.26e-3) 2.81e-2 (6.42e-3) 8.85e-2 (4.03e-2) 8.79e-2 (1.06e-1) 3.85e-1 (7.96e-1)
0.7 1.53e-1 (4.62e-2) 6.26e-2 (6.76e-2) 3.22e-2 (6.15e-3) 2.90e-2 (5.20e-3) 7.65e-2 (2.59e-2) 7.81e-2 (9.37e-2) 1.79e-1 (3.63e-1)
0.8 1.33e-1 (3.71e-2) 5.32e-2 (5.26e-2) 3.94e-2 (7.68e-3) 3.59e-2 (7.19e-3) 9.22e-2 (3.71e-2) 8.08e-2 (9.95e-2) 1.97e-1 (4.97e-1)
0.9 1.22e-1 (3.98e-2) 4.52e-2 (1.51e-2) 4.10e-2 (7.24e-3) 4.08e-2 (1.56e-2) 8.80e-2 (2.09e-2) 9.06e-2 (1.06e-1) 1.76e-1 (4.46e-1)
1.0 1.07e-1 (1.71e-2) 4.45e-2 (1.32e-2) 4.51e-2 (1.97e-2) 4.06e-2 (1.01e-2) 8.49e-2 (1.40e-2) 9.21e-2 (1.24e-1) 1.32e-1 (1.00e-1)
Table 8: Support recovery metrics for $p=128$, $s=10$, $\sigma=10^{-2}$.
$n/p$ Metric Penalty
$\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 TPR 4.81e-1 (1.51e-1) 2.72e-1 (1.37e-1) 2.38e-1 (1.54e-1) 2.40e-1 (1.63e-1) 1.70e-1 (1.30e-1) 9.70e-2 (9.37e-2) 1.81e-1 (1.24e-1)
FPR 4.68e-1 (4.63e-2) 2.30e-1 (5.82e-2) 1.95e-1 (6.49e-2) 1.96e-1 (8.75e-2) 1.39e-1 (3.23e-2) 5.77e-2 (1.92e-2) 1.58e-1 (3.41e-2)
F1 1.37e-1 (4.13e-2) 1.35e-1 (6.24e-2) 1.30e-1 (7.49e-2) 1.30e-1 (7.86e-2) 1.20e-1 (9.07e-2) 1.07e-1 (1.01e-1) 1.19e-1 (7.89e-2)
0.2 TPR 5.99e-1 (1.61e-1) 4.15e-1 (1.51e-1) 3.32e-1 (1.78e-1) 3.28e-1 (1.78e-1) 2.70e-1 (2.34e-1) 1.94e-1 (1.44e-1) 2.81e-1 (1.74e-1)
FPR 5.27e-1 (4.73e-2) 2.97e-1 (5.02e-2) 2.24e-1 (7.13e-2) 2.21e-1 (8.08e-2) 1.38e-1 (4.74e-2) 8.72e-2 (2.79e-2) 1.89e-1 (5.06e-2)
F1 1.53e-1 (4.02e-2) 1.69e-1 (5.94e-2) 1.66e-1 (8.15e-2) 1.64e-1 (7.72e-2) 1.91e-1 (1.70e-1) 1.79e-1 (1.44e-1) 1.64e-1 (1.20e-1)
0.3 TPR 7.13e-1 (1.66e-1) 5.24e-1 (1.69e-1) 5.03e-1 (2.41e-1) 4.77e-1 (2.14e-1) 5.80e-1 (4.06e-1) 4.35e-1 (3.11e-1) 5.23e-1 (3.39e-1)
FPR 5.29e-1 (5.16e-2) 2.97e-1 (7.97e-2) 2.26e-1 (9.84e-2) 2.15e-1 (1.08e-1) 1.04e-1 (4.23e-2) 6.98e-2 (5.77e-2) 1.43e-1 (7.10e-2)
F1 1.80e-1 (4.28e-2) 2.25e-1 (1.34e-1) 2.61e-1 (1.60e-1) 2.65e-1 (1.67e-1) 4.14e-1 (2.90e-1) 4.36e-1 (3.54e-1) 3.63e-1 (2.85e-1)
0.4 TPR 8.44e-1 (2.07e-1) 6.40e-1 (2.28e-1) 7.26e-1 (3.05e-1) 6.76e-1 (3.00e-1) 7.78e-1 (3.63e-1) 6.76e-1 (3.54e-1) 7.45e-1 (3.52e-1)
FPR 5.06e-1 (7.70e-2) 2.30e-1 (1.38e-1) 1.57e-1 (1.00e-1) 1.28e-1 (1.16e-1) 8.37e-2 (4.19e-2) 3.86e-2 (5.55e-2) 1.01e-1 (7.42e-2)
F1 2.19e-1 (6.38e-2) 3.99e-1 (3.36e-1) 4.49e-1 (2.43e-1) 4.97e-1 (3.03e-1) 5.71e-1 (2.73e-1) 6.83e-1 (3.80e-1) 5.58e-1 (3.06e-1)
0.5 TPR 9.20e-1 (1.91e-1) 8.07e-1 (2.46e-1) 8.56e-1 (2.83e-1) 8.60e-1 (2.79e-1) 9.01e-1 (2.63e-1) 8.00e-1 (3.05e-1) 8.49e-1 (2.94e-1)
FPR 3.94e-1 (9.21e-2) 1.20e-1 (1.64e-1) 1.15e-1 (8.48e-2) 7.77e-2 (1.02e-1) 7.51e-2 (3.57e-2) 2.85e-2 (6.12e-2) 8.36e-2 (6.74e-2)
F1 2.90e-1 (8.12e-2) 6.99e-1 (3.85e-1) 5.77e-1 (2.35e-1) 7.02e-1 (2.88e-1) 6.55e-1 (2.05e-1) 8.08e-1 (3.31e-1) 6.43e-1 (2.61e-1)
0.6 TPR 9.64e-1 (1.28e-1) 8.96e-1 (1.73e-1) 9.40e-1 (2.00e-1) 9.38e-1 (1.87e-1) 9.54e-1 (1.87e-1) 9.21e-1 (2.20e-1) 9.35e-1 (2.12e-1)
FPR 3.40e-1 (6.51e-2) 7.34e-2 (1.49e-1) 8.48e-2 (5.25e-2) 5.31e-2 (7.02e-2) 6.84e-2 (2.72e-2) 1.03e-2 (3.80e-2) 6.10e-2 (4.33e-2)
F1 3.29e-1 (6.20e-2) 8.28e-1 (3.16e-1) 6.62e-1 (1.74e-1) 7.81e-1 (2.11e-1) 6.96e-1 (1.53e-1) 9.26e-1 (2.28e-1) 7.26e-1 (1.92e-1)
0.7 TPR 9.93e-1 (7.00e-2) 9.52e-1 (1.09e-1) 9.94e-1 (4.45e-2) 9.81e-1 (1.00e-1) 9.92e-1 (8.00e-2) 9.60e-1 (1.46e-1) 9.95e-1 (5.00e-2)
FPR 3.23e-1 (4.41e-2) 2.86e-2 (9.83e-2) 7.13e-2 (3.38e-2) 3.28e-2 (3.16e-2) 6.13e-2 (2.03e-2) 3.81e-3 (2.23e-2) 4.86e-2 (2.40e-2)
F1 3.45e-1 (3.94e-2) 9.27e-1 (2.11e-1) 7.12e-1 (9.02e-2) 8.41e-1 (1.22e-1) 7.36e-1 (8.77e-2) 9.65e-1 (1.46e-1) 7.83e-1 (8.78e-2)
0.8 TPR 9.92e-1 (8.00e-2) 9.74e-1 (9.06e-2) 9.94e-1 (6.00e-2) 9.88e-1 (8.44e-2) 9.90e-1 (1.00e-1) 9.82e-1 (1.13e-1) 9.87e-1 (9.17e-2)
FPR 3.15e-1 (3.84e-2) 8.64e-3 (6.10e-2) 6.77e-2 (3.54e-2) 3.29e-2 (4.57e-2) 6.11e-2 (2.04e-2) 3.73e-3 (2.65e-2) 4.99e-2 (2.89e-2)
F1 3.50e-1 (4.00e-2) 9.75e-1 (1.21e-1) 7.24e-1 (9.11e-2) 8.54e-1 (1.19e-1) 7.35e-1 (9.71e-2) 9.81e-1 (1.24e-1) 7.77e-1 (1.08e-1)
0.9 TPR 9.95e-1 (5.00e-2) 9.83e-1 (7.66e-2) 9.96e-1 (4.00e-2) 9.90e-1 (1.00e-1) 1.00 (0.00) 9.90e-1 (8.10e-2) 9.92e-1 (8.00e-2)
FPR 3.20e-1 (4.52e-2) 3.73e-3 (3.73e-2) 6.59e-2 (2.71e-2) 2.57e-2 (1.44e-2) 6.00e-2 (2.02e-2) 1.44e-3 (1.44e-2) 4.77e-2 (2.41e-2)
F1 3.48e-1 (3.70e-2) 9.86e-1 (9.07e-2) 7.27e-1 (8.64e-2) 8.63e-1 (1.08e-1) 7.44e-1 (6.55e-2) 9.90e-1 (8.64e-2) 7.85e-1 (9.89e-2)
1.0 TPR 1.00 (0.00) 9.88e-1 (3.84e-2) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 9.99e-1 (1.00e-2) 1.00 (0.00)
FPR 3.19e-1 (3.86e-2) 0.00 (0.00) 7.05e-2 (2.47e-2) 2.97e-2 (1.52e-2) 6.36e-2 (2.35e-2) 0.00 (0.00) 4.95e-2 (2.07e-2)
F1 3.50e-1 (2.79e-2) 9.94e-1 (2.08e-2) 7.14e-1 (7.55e-2) 8.56e-1 (6.40e-2) 7.35e-1 (7.50e-2) 9.99e-1 (5.26e-3) 7.81e-1 (7.28e-2)
Table 9: Relative bias for large coefficients and estimated sparsity for $p=128$, $s=10$, $\sigma=10^{-2}$ (true sparsity = 10). Sparsity values are shown as mean (SD) in decimal format.
$n/p$ Metric Penalty
$\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 Bias 9.10e-1 (9.78e-2) 9.10e-1 (9.35e-2) 9.17e-1 (9.36e-2) 9.18e-1 (9.20e-2) 9.31e-1 (9.25e-2) 9.27e-1 (9.81e-2) 9.17e-1 (9.76e-2)
Sparsity 60.08 (5.81) 29.84 (7.37) 25.33 (8.43) 25.58 (11.25) 18.04 (3.90) 7.78 (2.30) 20.41 (4.12)
0.2 Bias 8.29e-1 (1.48e-1) 8.29e-1 (1.37e-1) 8.46e-1 (1.57e-1) 8.44e-1 (1.49e-1) 8.08e-1 (2.63e-1) 8.33e-1 (1.89e-1) 8.30e-1 (1.88e-1)
Sparsity 68.13 (5.74) 39.15 (6.17) 29.78 (9.16) 29.33 (10.50) 19.00 (5.09) 12.23 (2.88) 25.07 (5.86)
0.3 Bias 6.12e-1 (3.01e-1) 6.73e-1 (2.54e-1) 6.31e-1 (3.12e-1) 6.62e-1 (2.74e-1) 5.57e-1 (1.09e+0) 5.30e-1 (4.14e-1) 5.45e-1 (3.94e-1)
Sparsity 69.52 (6.05) 40.24 (9.09) 31.66 (11.34) 30.13 (12.62) 18.05 (3.30) 12.58 (4.44) 22.15 (5.90)
0.4 Bias 3.39e-1 (3.81e-1) 4.66e-1 (3.39e-1) 3.53e-1 (3.82e-1) 3.90e-1 (3.75e-1) 2.45e-1 (4.00e-1) 2.84e-1 (3.95e-1) 2.84e-1 (3.90e-1)
Sparsity 68.19 (8.60) 33.49 (14.47) 25.78 (10.35) 21.85 (12.22) 17.66 (2.99) 11.31 (3.54) 19.33 (5.85)
0.5 Bias 1.66e-1 (3.48e-1) 2.58e-1 (3.70e-1) 2.03e-1 (3.84e-1) 1.89e-1 (3.68e-1) 1.19e-1 (3.11e-1) 1.72e-1 (3.49e-1) 1.82e-1 (3.50e-1)
Sparsity 55.67 (10.06) 22.26 (17.17) 22.15 (8.19) 17.77 (10.07) 17.87 (2.87) 11.36 (4.70) 18.36 (5.36)
0.6 Bias 6.62e-2 (2.26e-1) 1.33e-1 (2.80e-1) 7.46e-2 (2.42e-1) 8.81e-2 (2.56e-1) 5.42e-2 (2.17e-1) 6.14e-2 (2.17e-1) 7.23e-2 (2.34e-1)
Sparsity 49.79 (6.98) 17.62 (16.08) 19.40 (4.85) 15.65 (6.77) 17.61 (2.96) 10.43 (2.54) 16.55 (3.46)
0.7 Bias 7.89e-3 (7.83e-2) 4.65e-2 (1.66e-1) 1.27e-2 (9.02e-2) 2.77e-2 (1.38e-1) 8.41e-3 (8.34e-2) 2.28e-2 (1.31e-1) 6.95e-3 (6.89e-2)
Sparsity 48.07 (5.02) 12.90 (10.74) 18.35 (3.77) 13.68 (3.35) 17.15 (2.41) 10.05 (1.52) 15.69 (2.55)
0.8 Bias 1.00e-2 (1.00e-1) 1.47e-2 (1.07e-1) 9.45e-3 (9.42e-2) 1.70e-2 (1.19e-1) 1.00e-2 (1.00e-1) 1.78e-2 (1.25e-1) 1.67e-2 (1.17e-1)
Sparsity 47.12 (4.38) 10.76 (6.51) 17.93 (3.73) 13.76 (4.60) 17.11 (2.37) 10.26 (2.02) 15.76 (2.79)
0.9 Bias 8.78e-3 (8.73e-2) 8.93e-3 (8.87e-2) 8.17e-3 (8.14e-2) 1.00e-2 (1.00e-1) 4.30e-5 (6.79e-5) 8.22e-3 (8.16e-2) 8.62e-3 (8.58e-2)
Sparsity 47.73 (5.31) 10.27 (3.73) 17.74 (3.02) 12.93 (2.12) 17.08 (2.38) 10.07 (0.91) 15.55 (2.47)
1.0 Bias 3.70e-5 (1.91e-5) 7.55e-5 (1.94e-4) 2.74e-5 (1.40e-5) 2.73e-5 (1.42e-5) 3.34e-5 (1.88e-5) 5.27e-5 (1.94e-4) 2.85e-5 (1.48e-5)
Sparsity 47.58 (4.55) 9.88 (0.38) 18.32 (2.91) 13.50 (1.80) 17.50 (2.78) 9.99 (0.10) 15.84 (2.44)
Table 10: RelErr for $p=128$, $s=10$, $\sigma=5\times 10^{-2}$.
$n/p$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 1.11 (1.03e-1) 1.12 (1.00e-1) 1.13 (1.09e-1) 1.12 (1.11e-1) 1.20 (1.25e-1) 1.21 (1.28e-1) 1.18 (1.11e-1)
0.2 1.05 (1.90e-1) 1.07 (1.60e-1) 1.04 (2.10e-1) 1.05 (1.86e-1) 1.04 (4.53e-1) 1.10 (3.09e-1) 1.11 (2.65e-1)
0.3 9.43e-1 (3.36e-1) 1.01 (2.85e-1) 8.95e-1 (3.46e-1) 9.16e-1 (3.08e-1) 7.15e-1 (6.20e-1) 8.65e-1 (5.28e-1) 9.36e-1 (5.18e-1)
0.4 4.95e-1 (5.28e-1) 7.43e-1 (4.81e-1) 4.60e-1 (5.14e-1) 5.65e-1 (4.83e-1) 3.60e-1 (5.59e-1) 4.56e-1 (5.77e-1) 4.92e-1 (6.01e-1)
0.5 2.14e-1 (4.45e-1) 4.04e-1 (5.17e-1) 2.29e-1 (4.22e-1) 2.87e-1 (4.33e-1) 1.42e-1 (3.88e-1) 2.00e-1 (4.37e-1) 2.53e-1 (4.84e-1)
0.6 1.00e-1 (3.43e-1) 1.67e-1 (3.94e-1) 9.37e-2 (3.07e-1) 1.12e-1 (3.31e-1) 8.69e-2 (3.19e-1) 9.24e-2 (3.15e-1) 1.03e-1 (3.34e-1)
0.7 1.38e-2 (1.32e-1) 3.40e-2 (1.92e-1) 1.14e-2 (1.11e-1) 2.33e-2 (1.62e-1) 2.29e-2 (1.58e-1) 1.20e-2 (1.05e-1) 2.54e-2 (1.76e-1)
0.8 5.71e-4 (2.14e-4) 9.45e-3 (8.94e-2) 2.96e-4 (1.25e-4) 2.84e-4 (1.23e-4) 4.45e-4 (1.75e-4) 8.97e-4 (2.06e-3) 3.76e-4 (1.52e-4)
0.9 4.92e-4 (1.63e-4) 5.58e-4 (8.14e-4) 2.56e-4 (1.06e-4) 2.42e-4 (1.03e-4) 3.83e-4 (1.45e-4) 5.53e-4 (3.29e-4) 3.22e-4 (1.26e-4)
1.0 4.55e-4 (1.37e-4) 3.43e-4 (4.74e-4) 2.43e-4 (7.81e-5) 2.29e-4 (7.53e-5) 3.46e-4 (1.22e-4) 4.31e-4 (2.00e-4) 2.89e-4 (9.78e-5)
Table 11: Computation time (seconds) for $p=128$, $s=10$, $\sigma=5\times 10^{-2}$.
$n/p$ $\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 9.64e-3 (3.51e-3) 2.44e-2 (1.48e-2) 5.18e-3 (3.54e-3) 4.40e-3 (1.78e-3) 1.68e-1 (5.43e-2) 7.64e-2 (1.91e-2) 1.23e-1 (6.09e-2)
0.2 2.24e-2 (7.28e-3) 5.94e-2 (2.20e-2) 4.74e-3 (1.77e-3) 4.76e-3 (8.76e-4) 2.13e-1 (5.68e-2) 9.69e-2 (1.59e-2) 2.48e-1 (9.15e-2)
0.3 6.26e-2 (1.97e-2) 7.69e-2 (3.31e-2) 6.34e-3 (3.71e-3) 6.24e-3 (3.23e-3) 1.43e-1 (7.94e-2) 9.20e-2 (3.46e-2) 4.88e-1 (2.73e-1)
0.4 2.27e-1 (6.77e-2) 1.14e-1 (7.67e-2) 2.13e-2 (1.13e-2) 1.78e-2 (9.01e-3) 1.63e-1 (1.46e-1) 1.63e-1 (1.24e-1) 7.96e-1 (7.34e-1)
0.5 2.82e-1 (3.66e-2) 1.12e-1 (8.95e-2) 2.72e-2 (8.83e-3) 2.41e-2 (7.18e-3) 1.25e-1 (1.09e-1) 1.41e-1 (1.34e-1) 6.28e-1 (8.97e-1)
0.6 2.07e-1 (6.77e-2) 7.37e-2 (7.15e-2) 2.91e-2 (6.88e-3) 2.66e-2 (5.30e-3) 1.02e-1 (8.82e-2) 9.80e-2 (9.68e-2) 3.73e-1 (7.51e-1)
0.7 1.49e-1 (4.21e-2) 4.52e-2 (3.38e-2) 2.90e-2 (6.20e-3) 2.62e-2 (4.91e-3) 7.64e-2 (3.72e-2) 7.39e-2 (7.60e-2) 1.81e-1 (4.57e-1)
0.8 1.43e-1 (3.61e-2) 4.35e-2 (1.32e-2) 4.00e-2 (4.28e-2) 3.27e-2 (8.47e-3) 8.86e-2 (1.76e-2) 8.16e-2 (6.47e-2) 1.23e-1 (4.85e-2)
0.9 1.19e-1 (2.18e-2) 4.13e-2 (1.17e-2) 3.71e-2 (7.24e-3) 3.37e-2 (5.48e-3) 8.54e-2 (1.52e-2) 7.90e-2 (4.17e-2) 1.14e-1 (2.01e-2)
1.0 1.10e-1 (1.90e-2) 3.99e-2 (8.52e-3) 3.95e-2 (7.56e-3) 3.68e-2 (6.30e-3) 8.91e-2 (2.66e-2) 7.06e-2 (3.83e-2) 1.17e-1 (3.60e-2)
Table 12: Support recovery metrics for $p=128$, $s=10$, $\sigma=5\times 10^{-2}$.
$n/p$ Metric Penalty
$\ell_{1}$ $\ell_{1/2}$ SCAD MCP TL1 LOG EXP
0.1 TPR 5.38e-1 (1.69e-1) 2.83e-1 (1.56e-1) 2.56e-1 (1.57e-1) 2.40e-1 (1.59e-1) 1.73e-1 (1.29e-1) 9.70e-2 (9.79e-2) 1.93e-1 (1.31e-1)
FPR 4.69e-1 (4.93e-2) 2.39e-1 (6.96e-2) 1.97e-1 (6.72e-2) 1.85e-1 (7.92e-2) 1.44e-1 (2.92e-2) 6.16e-2 (1.69e-2) 1.56e-1 (3.30e-2)
F1 1.53e-1 (4.76e-2) 1.39e-1 (7.32e-2) 1.41e-1 (7.85e-2) 1.37e-1 (8.21e-2) 1.18e-1 (8.67e-2) 1.03e-1 (9.33e-2) 1.26e-1 (7.98e-2)
0.2 TPR 6.14e-1 (1.56e-1) 4.34e-1 (1.48e-1) 3.54e-1 (1.65e-1) 3.45e-1 (1.71e-1) 3.04e-1 (2.67e-1) 2.17e-1 (1.44e-1) 3.11e-1 (1.69e-1)
FPR 5.30e-1 (5.00e-2) 3.02e-1 (5.51e-2) 2.31e-1 (7.14e-2) 2.27e-1 (9.19e-2) 1.37e-1 (4.88e-2) 8.67e-2 (3.00e-2) 1.86e-1 (4.73e-2)
F1 1.57e-1 (4.06e-2) 1.76e-1 (6.51e-2) 1.75e-1 (8.32e-2) 1.73e-1 (8.37e-2) 2.15e-1 (2.04e-1) 2.03e-1 (1.59e-1) 1.83e-1 (1.19e-1)
0.3 TPR 6.62e-1 (1.74e-1) 5.09e-1 (1.78e-1) 4.68e-1 (2.38e-1) 4.43e-1 (2.23e-1) 5.35e-1 (4.10e-1) 3.76e-1 (2.98e-1) 4.59e-1 (3.08e-1)
FPR 5.34e-1 (4.97e-2) 2.90e-1 (8.03e-2) 2.20e-1 (9.78e-2) 2.00e-1 (1.06e-1) 1.09e-1 (4.63e-2) 7.95e-2 (5.31e-2) 1.63e-1 (6.64e-2)
F1 1.67e-1 (4.40e-2) 2.23e-1 (1.36e-1) 2.42e-1 (1.46e-1) 2.50e-1 (1.60e-1) 3.78e-1 (2.92e-1) 3.64e-1 (3.33e-1) 2.98e-1 (2.47e-1)
0.4 TPR 8.31e-1 (2.07e-1) 6.26e-1 (2.46e-1) 7.11e-1 (3.35e-1) 6.48e-1 (3.03e-1) 7.52e-1 (3.81e-1) 6.47e-1 (3.56e-1) 7.04e-1 (3.64e-1)
FPR 4.90e-1 (8.25e-2) 2.32e-1 (1.43e-1) 1.54e-1 (9.70e-2) 1.44e-1 (1.16e-1) 8.46e-2 (3.78e-2) 4.57e-2 (6.17e-2) 1.06e-1 (7.45e-2)
[…]